# PIXEL: Adaptive Steering Via Position-wise Injection with eXact Estimated Levels under Subspace Calibration

### v1

@@ 1 Introduction
Large Language Models (LLMs) have achieved remarkable success across a broad spectrum of natural language processing tasks, ranging from question answering to open-ended generation (Brown et al., 2020; Touvron et al., 2023; OpenAI, 2023). However, ensuring that LLMs behave in alignment with desired attributes such as truthfulness, fairness, or informativeness remains an open challenge, especially in real-world deployments where outputs must satisfy domain-specific or ethical constraints (Ouyang et al., 2022; Bai et al., 2022).
To meet this need, a growing body of work has focused on post-training control mechanisms, which aim to adjust model behavior without retraining the entire model. Among them, activation steering has emerged as a promising lightweight approach that manipulates internal representations at inference time to guide outputs toward target properties (Rimsky et al., 2024a; Li et al., 2023b; Im & Li, 2025). Compared to supervised fine-tuning (Wei et al., 2022)(SFT) or reinforcement learning with human feedback (Christiano et al., 2017)(RLHF), steering methods require no additional gradient updates and can be flexibly applied to frozen LLM. Most existing methods construct steering based on the activation difference between positive and negative samples and apply a shift with a strength controlled by a coefficient at selected layers or locations.
@@
Recent work has introduced adaptability into activation control. ACT  (Wang et al., 2024)trains probes and clusters direction vectors into categories, then scales the edits based on the probe score of the selected head, adjusting the strength of the intervention at inference. SADI  (Zhang et al., 2025) dynamically adjusts the control direction by performing element-wise edits on the connection activations of all layers, further adjusting the intervention based on the input semantics. However, neither approach provides a principled treatment of the location and extent of the intervention. ACT relies on learned detectors and task-tuned hyperparameters, which limit its robustness and transferability across attribute-aligned datasets. While SADI does not share ACT’s limitations, it requires uniform application of interventions across all layers, preventing adaptive selection of the location of injected steering vectors.
@@ 2.1 Activation Steering Methods
Activation steering refers to modifying hidden representations of large language models at inference time to adjust outputs toward desired attributes, without updating model parameters (Im & Li, 2025; Cao et al., 2024; Bayat et al., 2025; Oozeer et al., 2025; Subramani et al., 2024; Liu et al., 2024b; Hernandez et al., 2023). A variety of methods have been proposed in this line of work. Contrastive Activation Addition (CAA) (Rimsky et al., 2024b) computes steering vectors by averaging activation differences between positive and negative examples and injects them at inference to shift model behavior. Inference-Time Intervention (ITI) (Li et al., 2023a) applies fixed steering directions across attention heads to enhance factuality and truthfulness. Representation Engineering (Olson et al., 2024) explores editing latent vectors to induce desired outputs, while Rank-One Model Editing (ROME) (Meng et al., 2022) performs low-rank updates to affect factual memory in LMs. AutoSteer (Yang et al., 2024) automates the discovery of steering directions by combining unsupervised clustering and contrastive signals.
@@ 2.2 Adaptive Control
Adaptive control seeks to dynamically adjust the strength or weighting of steering at inference time, rather than relying on fixed heuristics or hand-crafted schedules. In this line of work, a range of approaches has been proposed in recent work (Ross et al., 2022; Geva et al., 2023; Lin et al., 2023; Li et al., 2023b; Wang et al., 2025; Sun et al., 2024; Subramani et al., 2024), which instantiate adaptivity through different signals (e.g., probe predictions, semantic cues, or contrastive objectives) and at different granularities (per layer/head/token). For example, Adaptive Activation Steering (ACT) (Wang et al., 2025) trains probes to detect steering patterns and adaptively scales intervention strength based on probe outputs. SADI (Zhang et al., 2025) introduces semantics-adaptive dynamically interventions, adjusting activation vectors according to input semantics, while FairSteer (Nguyen et al., 2025a) detects biased representations during inference and applies debiasing steering vectors only when necessary. Contrastive Null Steer (Sun et al., 2024) further improves adaptivity by learning to steer away from undesired directions via contrastive supervision. Token-wise Adaptive Modulation (TAM) (Liu et al., 2024a) modulates intervention strength at each token position using auxiliary gating networks, enabling fine-grained control over steering dynamics.
@@ 3.3 Representation-Level Guarantees
$\Big|\;\mathbb{E}[\tilde{\phi}(x)]\;-\;\tfrac{1}{n}\sum_{i=1}^{n}\tilde{\phi}(x_{i})\;\Big|\;\leq\;\sqrt{\tfrac{\ln(2/\delta)}{2n}}.$
@@
Assume $\mathcal{S}\neq\varnothing$ and for each $\ell\in\mathcal{S}$ the strength $\alpha^{*}_{\ell,t^{\star}}$ is chosen as the minimal nonnegative value such that $\cos\!\big(h^{\prime}_{\ell,t^{\star}}(x),\,w_{\ell}(x)\big)\geq s$. Then for every sample $x$ and every $\ell\in\mathcal{S}$,
$\phi_{\mathrm{site}}^{(\ell)}(x)=\cos(h^{\prime}_{\ell,t^{\star}}(x),w_{\ell}(x))-s\;\geq\;0,\quad\text{and hence}\quad\bar{\phi}_{\mathrm{margin}}(x)\;\geq\;0.$
Remark At each selected site we decompose the representation into a component parallel to $w_{\ell}$ and an orthogonal remainder, then add the smallest nonnegative increment along $w_{\ell}$ so that the cosine meets the target $s$ (leaving the orthogonal part unchanged). By monotonicity the site-wise—and hence averaged—margin is nonnegative; the increment is zero iff the original cosine already meets $s$ and, when needed, increases strictly with $s$ (see Lemma 3, 2).
@@ 4.1 Experimental Settings
4.1.1 Datasets and Metrics.
We evaluate PIXEL on four widely used alignment suites, each with its standard metric and judging protocol:
@@
4.1.2 Models and Baselines.
We consider three representative large language models for experiments:
@@
  • ITI (Li et al., 2023a): inference-time intervention that perturbs hidden states to reshape the output distribution.
  • CAA (Rimsky et al., 2024b): contrastive activation addition using differences between positive/negative exemplars.
@@
4.1.3 Experimental Setup.
All experiments were conducted on NVIDIA RTX 6000 Ada GPUs. For all datasets, subspace extraction used 200 samples. We used subspace similarity thresholds of 0.85 and 0.9. For each configuration, we report the average and standard deviation over three runs with different random seeds.
@@ 4.2 Main results
4.2.1 Comprehensive gains across heterogeneous evaluation forms.
@@
PIXEL achieves significant improvements under two evaluation paradigms: multiple-choice classification (TruthfulQA MC1/MC2 for factuality; BBQ accuracy for bias) and open-ended generation (auto judges score the rejection of unsafe requests; HelpSteer-Help scores helpfulness on a scale of 0–4). Taking Qwen2-7B as the representative base, PIXEL achieves 37.50/55.22 on MC1 and MC2, 0.732 on BBQ, 0.66 on Rejection, and 3.94 on HelpSteer-Help; these correspond to relative gains of 42%, 22%, 15%, 74%, and 12% relative to the base model (26.38/45.41, 0.634, 0.38, 3.51), and improvements of 26%, 13%, 15%, 57%, and 6% relative to the strongest competing baseline in this setting (29.83/48.69, 0.636, 0.42, 3.73). We attribute these broad gains to an answer-agnostic approach to constructing guidance: rather than forming pairs by associating known correct and incorrect answers (an approach that entangles the learned signal with format-specific templates and labels), we incorporate generic attribute-related instructions (e.g., Avoid repeating myths; respond concisely with verified facts.) with the original question (e.g., What happens if you prick your finger on a spindle?). This decoupling enables the learned directions to transfer from truthfulness and bias detection in multiple-choice questions to rejection and assistance in open-ended generation. Furthermore, dual-view refinement within a layer (tail-window averaging and end-token differentials) provides complementary information, capturing both stable conditioning shifts and localized changes relevant to initial decoding, thereby supporting comprehensive improvements across evaluation modalities.
4.2.2 PIXEL maintains model performance on general capability benchmarks.
@@
To ensure that our steering approach does not compromise the model’s inherent language understanding and reasoning capabilities, we evaluate its performance on a range of standard NLP benchmarks, including RACE (reading comprehension), MMLU (multi-task knowledge), OpenBookQA (commonsense reasoning), and GLUE (general language understanding). The results in Table 2 show that PIXEL not only maintains the performance of the base model but also improves it in some cases, highlighting its precision in attribute alignment without compromising core performance. This stands in stark contrast to baseline methods, which often suffer from performance trade-offs. For example, while CAA relies on explicit positive/negative examples and excels on knowledge-focused closed-set tasks such as MMLU and GLUE (Qwen2 scores 0.830 on GLUE), it exhibits instability on complex reasoning tasks (0.625 on RACE), suggesting that its steering vectors may overfit to simple fact contrasts at the expense of nuanced reasoning paths. ReFT improves one metric (0.677 on RACE for Llama3) but underperforms another (MMLU falls below 0.651 on the original model), indicating limited generalization. Most notably, ITI’s fixed-weight intervention leads to a significant drop in knowledge-intensive benchmarks (MMLU drops to 0.546 on Llama3), likely because the coarse-grained activation steering undermines the model’s generalizability. In contrast, PIXEL performs position-adaptive steering, applying minimal geometric consistency intervention only at the most acceptable locations within the model’s hidden representation. This design ensures minimal disruption to the underlying model’s behavior. Intuitively, if alignment with the target attribute can be achieved with a smaller steering strength adjustment, it may interfere less with the model’s capabilities on other attributes than achieving the same effect with a larger steering strength.
@@ 4.3 Ablation Studies
Figure 3: Analysis of $z_{\text{target}}$ sensitivity in PIXEL.
@@
4.3.1 Validating the Effectiveness of the Proposed Dual-View Property-Aligned Subspace
We designed a PIXEL variant (Single-View) that removes the tail average view and uses only the last token difference to construct the attribute subspace. Figure 2 compares the performance of Single-View and Dual-View on various datasets. As shown, Dual-View improves TruthfulQA performance across all backbone networks. On MC1, Mistral’s score improves from 31.2 to 34.1 (+2.9), Qwen’s score improves from 30.5 to 36.4 (+5.9), and Llama’s score improves from 32.0 to 46.8 (+14.8). MC2 also showed similar improvements: Mistral’s score increased from 52.1 to 54.3 (+2.2), Qwen’s score increased from 55.0 to 59.2 (+4.2), and Llama’s score increased from 51.3 to 60.5 (+9.2). Dual-view also outperformed Single-View on BBQ (accuracy) across all three backbones. These results suggest that restricting the subspace to terminal label signals underutilizes cues distributed along the tail of the response, thereby limiting attribute alignment for factuality indicators. By aggregating complementary views, the dual-view construction provides a more robust, attribute-aligned global subspace, making the model’s activation steering more accurate and stable.
4.3.2 Evaluating the effectiveness of the adaptive steering strength mechanism.
To evaluate the impact of steering strength regulation, we designed a PIXEL variant(w/o adaptive ) that disables the adaptive weighting mechanism and instead applies a fixed intervention magnitude $\alpha=1$ across all positions and layers. Table 3 shows a comparison between this fixed-strength variant and the full PIXEL. Across all models and evaluation settings, the adaptive strategy consistently leads to stronger performance. For example, on Llama3-8B-Instruct, using adaptive $\alpha$ yields an MC2 score of 60.28, compared to 51.20 without it; on Qwen2-7B-Instruct, the rejection rate improves from 0.42 to 0.66. These results highlight two core limitations of fixed-strength intervention. First, a uniform intensity can be insufficient at locations where critical steering is required, leading to understeering and limited behavioral change. Second, indiscriminately applying the same magnitude can induce oversteering at places that are semantically inconsistent with the target orientation, leading to poor generalization or unnatural generation. In contrast, PIXEL utilizes closed-form optimization to compute the minimum required steering magnitude at each location, subject to a cosine similarity threshold relative to the target subspace. This adaptive mechanism enables the model to inject only when necessary and with just enough strength to achieve alignment, balancing effectiveness and safety. The observed improvements demonstrate that precise, geometry-aware interventions are more robust and transferable than coarse, global interventions.
@@ 4.3.1 Validating the Effectiveness of the Proposed Dual-View Property-Aligned Subspace
4.3.3 Validating the effectiveness of the proposed orthogonal residual calibration mechanism.
To evaluate the impact of orthogonal residual calibration, we designed a PIXEL variant (w/o calibration) that disables per-sample orthogonal residuals and adjusts only along global, attribute-aligned directions. Table 4 compares this variant with PIXEL. Disabling calibration leads to consistent performance degradation across models and datasets. For example, on Mistral-7B-v0.3, the MC1 score drops from 33.80 to 29.17 (-13.7%); on Qwen2-7B-Instruct, the usefulness drops from 3.94 to 3.51 (-11.0%); and on Llama3-8B-Instruct, the MC1 drops from 46.30 to 43.06 (-7.0%). These results demonstrate two limitations of using only the global subspace without calibration. First, it cannot adapt to input-specific semantic shifts, so steering may miss the subspace most relevant to a given cue, suppressing gains in factuality, bias, and safety. Second, it suffers from volatility in open-ended environments: without calibration, interventions may push generations into regions of poor alignment, which manifest as decreased helpfulness or inconsistent rejection. By introducing orthogonal residual calibration, alignment can be improved without changing the injection strength, resulting in more robust and transferable gains.

### v2

@@ 1 Introduction
Large Language Models (LLMs) (Brown et al., 2020; Touvron et al., 2023; OpenAI, 2023) have achieved remarkable success across a broad spectrum of natural language processing tasks, ranging from question answering to open-ended generation Su et al. (2023b; a); Cheng et al. (2024); Yang et al. (2024a); Zhang et al. (2024); Yang et al. (2025a); Hu et al. (2024a). However, ensuring that LLMs behave in alignment with desired attributes such as truthfulness, fairness, or informativeness remains an open challenge, especially in real-world deployments where outputs must satisfy domain-specific or ethical constraints (Ouyang et al., 2022; Bai et al., 2022; Yang et al., 2024b; Zhang et al., 2025b; Wang et al., 2025a; Zhou et al., 2025; Dong et al., 2025; Yang et al., 2025c; Guo et al., 2025).
To meet this need, a growing body of work has focused on post-training control mechanisms, which aim to adjust model behavior without retraining the entire model. Among them, activation steering has emerged as a promising lightweight approach that manipulates internal representations at inference time to guide outputs toward target properties (Rimsky et al., 2024b; Li et al., 2023a; Yang et al., 2024d; Hu et al., 2024b). Compared to supervised fine-tuning (Wei et al., 2022)(SFT) or reinforcement learning with human feedback (Christiano et al., 2017)(RLHF), steering methods require no additional gradient updates and can be flexibly applied to frozen LLM. Most existing methods construct steering based on the activation difference between positive and negative samples and apply a shift with a strength controlled by a coefficient at selected layers or locations.
@@
Recent work has introduced adaptability into activation control. ACT  (Wang et al., 2024)trains probes and clusters direction vectors into categories, then scales the edits based on the probe score of the selected head, adjusting the strength of the intervention at inference. SADI  (Zhang et al., 2025a) dynamically adjusts the control direction by performing element-wise edits on the connection activations of all layers, further adjusting the intervention based on the input semantics. However, neither approach provides a principled treatment of the location and extent of the intervention. ACT relies on learned detectors and task-tuned hyperparameters, which limit its robustness and transferability across attribute-aligned datasets. While SADI does not share ACT’s limitations, it requires uniform application of interventions across all layers, preventing adaptive selection of the location of injected steering vectors.
@@ 2.1 Activation Steering Methods
Activation steering refers to modifying hidden representations of large language models at inference time to adjust outputs toward desired attributes, without updating model parameters (Im and Li, 2025; Cao et al., 2024; Bayat et al., 2025; Oozeer et al., 2025; Subramani et al., 2024; Liu et al., 2024a; Hernandez et al., 2023; Yang et al., 2025b; Yao et al., 2025; Jiang et al., 2025). A variety of methods have been proposed in this line of work. Contrastive Activation Addition (CAA) (Rimsky et al., 2024a) computes steering vectors by averaging activation differences between positive and negative examples and injects them at inference to shift model behavior. Inference-Time Intervention (ITI) (Li et al., 2023b) applies fixed steering directions across attention heads to enhance factuality and truthfulness. Representation Engineering (Olson et al., 2024) explores editing latent vectors to induce desired outputs, while Rank-One Model Editing (ROME) (Meng et al., 2022) performs low-rank updates to affect factual memory in LMs. AutoSteer (Yang et al., 2024c) automates the discovery of steering directions by combining unsupervised clustering and contrastive signals.
@@ 2.2 Adaptive Control
Adaptive control seeks to dynamically adjust the strength or weighting of steering at inference time, rather than relying on fixed heuristics or hand-crafted schedules. In this line of work, A range of approaches have been proposed in recent work (Ross et al., 2022; Geva et al., 2023; Lin et al., 2023; Li et al., 2023a; Wang et al., 2025b; Sun et al., 2024; Subramani et al., 2024), which instantiate adaptivity through different signals (e.g., probe predictions, semantic cues, or contrastive objectives) and at different granularities (per layer/head/token). For example, Adaptive Activation Steering (ACT) (Wang et al., 2025b) trains probes to detect steering patterns and adaptively scales intervention strength based on probe outputs. SADI (Zhang et al., 2025a) introduces semantics-adaptive dynamically interventions, adjusting activation vectors according to input semantics, while FairSteer (Nguyen et al., 2025a) detects biased representations during inference and applies debiasing steering vectors only when necessary. Contrastive Null Steer (Sun et al., 2024) further improves adaptivity by learning to steer away from undesired directions via contrastive supervision. Token-wise Adaptive Modulation (TAM) (Liu et al., 2024b) modulates intervention strength at each token position using auxiliary gating networks, enabling fine-grained control over steering dynamics.
@@ 3.3 Representation-Level Guarantees
$\Bigl|\;\mathbb{E}[\tilde{\phi}(x)]\;-\;\tfrac{1}{n}\sum_{i=1}^{n}\tilde{\phi}(x_{i})\;\Bigr|\;\leq\;\sqrt{\tfrac{\ln(2/\delta)}{2n}}.$
@@
Assume $\mathcal{S}\neq\varnothing$ and for each $\ell\in\mathcal{S}$ the strength $\alpha^{*}_{\ell,t^{\star}}$ is chosen as the minimal nonnegative value such that $\cos\!\bigl(h^{\prime}_{\ell,t^{\star}}(x),\,w_{\ell}(x)\bigr)\geq s$. Then for every sample $x$ and every $\ell\in\mathcal{S}$,
$\phi_{\mathrm{site}}^{(\ell)}(x)=\cos\!\bigl(h^{\prime}_{\ell,t^{\star}}(x),\,w_{\ell}(x)\bigr)-s\;\geq\;0,\quad\text{and hence}\quad\bar{\phi}_{\mathrm{margin}}(x)\;\geq\;0.$
Remark At each selected site we decompose the representation into a component parallel to $w_{\ell}$ and an orthogonal remainder, then add the smallest nonnegative increment along $w_{\ell}$ so that the cosine meets the target $s$ (leaving the orthogonal part unchanged). By monotonicity the site-wise—and hence averaged—margin is nonnegative; the increment is zero iff the original cosine already meets $s$ and, when needed, increases strictly with $s$ (see Lemma 3 and Lemma 2).
@@ 4.1 Experimental Settings
Datasets and Metrics. We evaluate PIXEL on four widely used alignment suites, each with its standard metric and judging protocol:
@@
Models and Baselines. We consider three representative large language models for experiments:
@@
  • ITI (Li et al., 2023b): inference-time intervention that perturbs hidden states to reshape the output distribution.
  • CAA (Rimsky et al., 2024a): contrastive activation addition using differences between positive/negative exemplars.
@@
Experimental Setup. All experiments were conducted on NVIDIA RTX 6000 Ada GPUs. For all datasets, subspace extraction used 200 samples. We used subspace similarity thresholds of 0.85 and 0.9. For each configuration, we report the average and standard deviation over three runs with different random seeds.
@@ 4.2 Main results
Comprehensive gains across heterogeneous evaluation forms. PIXEL achieves significant improvements under two evaluation paradigms: multiple-choice classification (TruthfulQA MC1/MC2 for factuality; BBQ accuracy for bias) and open-ended generation (auto judges score the rejection of unsafe requests; HelpSteer-Help scores helpfulness on a scale of 0–4). Taking Qwen2-7B as the representative base, PIXEL achieves 37.50/55.22 on MC1 and MC2, 0.732 on BBQ, 0.66 on Rejection, and 3.94 on HelpSteer-Help; these correspond to relative gains of 42%, 22%, 15%, 74%, and 12% relative to the base model (26.38/45.41, 0.634, 0.38, 3.51), and improvements of 26%, 13%, 15%, 57%, and 6% relative to the strongest competing baseline in this setting (29.83/48.69, 0.636, 0.42, 3.73). We attribute these broad gains to an answer-agnostic approach to constructing guidance: rather than forming pairs by associating known correct and incorrect answers (an approach that entangles the learned signal with format-specific templates and labels), we incorporate generic attribute-related instructions (e.g., Avoid repeating myths; respond concisely with verified facts.) with the original question (e.g., What happens if you prick your finger on a spindle?). This decoupling enables the learned directions to transfer from truthfulness and bias detection in multiple-choice questions to rejection and assistance in open-ended generation. Furthermore, dual-view refinement within a layer (tail-window averaging and end-token differentials) provides complementary information, capturing both stable conditioning shifts and localized changes relevant to initial decoding, thereby supporting comprehensive improvements across evaluation modalities.
@@
PIXEL maintains model performance on general capability benchmarks. To ensure that our steering approach does not compromise the model’s inherent language understanding and reasoning capabilities, we evaluate its performance on a range of standard NLP benchmarks, including RACE (reading comprehension), MMLU (multi-task knowledge), OpenBookQA (commonsense reasoning), and GLUE (general language understanding). The results in Table 2 show that PIXEL not only maintains the performance of the base model but also improves it in some cases, highlighting its precision in attribute alignment without compromising core performance. This stands in stark contrast to baseline methods, which often suffer from performance trade-offs. For example, while CAA relies on explicit positive/negative examples and excels on knowledge-focused closed-set tasks such as MMLU and GLUE (Qwen2 scores 0.830 on GLUE), it exhibits instability on complex reasoning tasks (0.625 on RACE), suggesting that its steering vectors may overfit to simple fact contrasts at the expense of nuanced reasoning paths. ReFT improves one metric (0.677 on RACE for Llama3) but underperforms another (MMLU falls below 0.651 on the original model), indicating limited generalization. Most notably, ITI’s fixed-weight intervention leads to a significant drop in knowledge-intensive benchmarks (MMLU drops to 0.546 on Llama3), likely because the coarse-grained activation steering undermines the model’s generalizability. In contrast, PIXEL performs position-adaptive steering, applying minimal geometric consistency intervention only at the most acceptable locations within the model’s hidden representation. This design ensures minimal disruption to the underlying model’s behavior. Intuitively, if alignment with the target attribute can be achieved with a smaller steering strength adjustment, it may interfere less with the model’s capabilities on other attributes than achieving the same effect with a larger steering strength.
@@ 4.3 Ablation Studies
Validating the Effectiveness of the Proposed Dual-View Property-Aligned Subspace. We designed a PIXEL variant (Single-View) that removes the tail average view and uses only the last token difference to construct the attribute subspace. Figure 2 compares the performance of Single-View and Dual-View on various datasets. As shown, Dual-View improves TruthfulQA performance across all backbone networks. On MC1, Mistral’s score improves from 31.2 to 34.1 (+2.9), Qwen’s score improves from 30.5 to 36.4 (+5.9), and Llama’s score improves from 32.0 to 46.8 (+14.8). MC2 also showed similar improvements: Mistral’s score increased from 52.1 to 54.3 (+2.2), Qwen’s score increased from 55.0 to 59.2 (+4.2), and Llama’s score increased from 51.3 to 60.5 (+9.2). Dual-view also outperformed Single-View on BBQ (accuracy) across all three backbones. These results suggest that restricting the subspace to terminal label signals underutilizes cues distributed along the tail of the response, thereby limiting attribute alignment for factuality indicators. By aggregating complementary views, the dual-view construction provides a more robust, attribute-aligned global subspace, making the model’s activation steering more accurate and stable.
Evaluating the effectiveness of the adaptive steering strength mechanism. To evaluate the impact of steering strength regulation, we designed a PIXEL variant(w/o adaptive ) that disables the adaptive weighting mechanism and instead applies a fixed intervention magnitude $\alpha=1$ across all positions and layers. Table 3 shows a comparison between this fixed-strength variant and the full PIXEL. Across all models and evaluation settings, the adaptive strategy consistently leads to stronger performance. For example, on Llama3-8B-Instruct, using adaptive $\alpha$ yields an MC2 score of 60.28, compared to 51.20 without it; on Qwen2-7B-Instruct, the rejection rate improves from 0.42 to 0.66. These results highlight two core limitations of fixed-strength intervention. First, a uniform intensity can be insufficient at locations where critical steering is required, leading to understeering and limited behavioral change. Second, indiscriminately applying the same magnitude can induce oversteering at places that are semantically inconsistent with the target orientation, leading to poor generalization or unnatural generation. In contrast, PIXEL utilizes closed-form optimization to compute the minimum required steering magnitude at each location, subject to a cosine similarity threshold relative to the target subspace. This adaptive mechanism enables the model to inject only when necessary and with just enough strength to achieve alignment, balancing effectiveness and safety. The observed improvements demonstrate that precise, geometry-aware interventions are more robust and transferable than coarse, global interventions.
@@
Validating the effectiveness of the proposed orthogonal residual calibration mechanism. To evaluate the impact of orthogonal residual calibration, we designed a PIXEL variant (w/o calibration) that disables per-sample orthogonal residuals and adjusts only along global, attribute-aligned directions. Table 4 compares this variant with PIXEL. Disabling calibration leads to consistent performance degradation across models and datasets. For example, on Mistral-7B-v0.3, the MC1 score drops from 33.80 to 29.17 (-13.7%); on Qwen2-7B-Instruct, the usefulness drops from 3.94 to 3.51 (-11.0%); and on Llama3-8B-Instruct, the MC1 drops from 46.30 to 43.06 (-7.0%). These results demonstrate two limitations of using only the global subspace without calibration. First, it cannot adapt to input-specific semantic shifts, so steering may miss the subspace most relevant to a given cue, suppressing gains in factuality, bias, and safety. Second, it suffers from volatility in open-ended environments: without calibration, interventions may push generations into regions of poor alignment, which manifest as decreased helpfulness or inconsistent rejection. By introducing orthogonal residual calibration, alignment can be improved without changing the injection strength, resulting in more robust and transferable gains.